OpenStudy (anonymous):

'Demuestra que' means 'show that'. I'm so confused. Please help me with this question.

OpenStudy (anonymous):

This is the question!

OpenStudy (john_es):

Use logarithms properties,. \[\log\left[(x+\sqrt{x^2-1})(x-\sqrt{x^2-1})\right]=\log\left[x^2-(x^2-1)\right]=\log1=0\]